A pizza shop offers the following toppings: 8 different vegetables, 5 different meats, and 4 different cheeses.
How many ways can 4 toppings be selected where:
a.All the toppings are vegetables
b.All the toppings are meat
c.There is only cheese on the pizza
d.There are 2 vegetables, 1 meat and 1 cheese
e.There are 3 meats and 1 cheese.
: A pizza shop offers the following toppings: 8 different vegetables, 5 different meats, and 4 different cheeses.
How many ways can 4 toppings be selected where:
a.All the toppings are vegetables
b.All the toppings are meat
c.There is only cheese on the pizza
d.There are 2 vegetables, 1 meat and 1 cheese
e.There are 3 meats and 1 cheese.

Answer by checkley75(3666)

8*7*6*5=1680 DIFFERENT TOPPINGS OF VEGETABLES.
5*4*3*2=120 MEAT TOPPINGS.
4*3*2*1=24 CHEEZE TOPPINGS.
8*7*5*4=1120 MIXED TOPPINGS.
5*4*3*4=240 MIXED MEAT & CHEEZE.